Shift report 20010812 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Aug 12 2001 - 07:43:12 EDT

  • Next message: Apache: "Shift report 20010812 08:00-16:00"

    ------------------------------------------------------------
    Date: Sun Aug 12 07:43:12 2001, Shift: 00:00-08:00
    Supervisor: Michael Murray, with: Mads Mikelsen
    Log book#7 pages 64 - 66
    
    Shift Summary:
    ------------------------------------------------------------
    No beam due to bad vacuum in our DX magnet which caused
    it to quench. Since about 4pm their have been power dips and
    a failure in the tandem.
    
    There was a problem with the Wave Form Generator (WFG)
    loopback measurement of D3. It seemed to be a factor of
    2 higer than the measured current. This was solved by a 
    software reset of our Front End Computer (FEC).
    
    I (Michael) could not measure a pedestal on the ZDCs even
    when I disconnected the input to the ADC.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Aug 12 2001 - 07:44:03 EDT